Imagine that you are in the market to purchase a new home. The sales representative informs you that there are 56 houses available with two differant floor plans. use x to represent floor plan 1 and y to represent floor plan 2 to create an equation that illustrates this situation if there are three times as many homes available with floor plan 2 than floor plan 1. I do not understand if I am doing it correctly when I make the equation
x+3y=56 ???

Answer by nerdybill(7384)

Imagine that you are in the market to purchase a new home. The sales representative informs you that there are 56 houses available with two differant floor plans. use x to represent floor plan 1 and y to represent floor plan 2 to create an equation that illustrates this situation if there are three times as many homes available with floor plan 2 than floor plan 1.
.
Let x = number of floor plan 1 homes
and y = number of floor plan 2 homes
then from: "there are 56 houses available" we get:
x + y = 56 (equation 1)
and from: "there are three times as many homes available with floor plan 2 than floor plan 1" we get:
y = 3x (equation 2)
.
So, using the definition from equation 2, we substitute into equation 1:
starting with equation 1:
x + y = 56
x + 3x = 56 (I think this is more of what they wanted)